# LLM Braces: Straightening Out LLM Predictions with Relevant Sub-Updates

### v1

@@ Abstract
Recent findings reveal that much of the knowledge in a Transformer-based Large Language Model (LLM) is encoded in its feed-forward (FFN) layers, where each FNN layer can be interpreted as the summation of sub-updates, each corresponding to a weighted column vector from the FFN’s value parameter matrix that often encodes human-interpretable concepts. In light of this, we hypothesize that model performance and behaviors can be further enhanced and controlled by modulating the contributions of these sub-updates based on their relevance to the input or target output style, and propose LLMBraces, a novel and efficient method that computes relevance scores associated with value vectors in FFN layers and leverages these scores to dynamically adjust the contribution of sub-updates. By optimizing sub-update contributions, LLMBraces refines the prediction process, leading to more accurate and reliable outputs, much like a ‘brace’ providing support and stability. Moreover, LLMBraces can be extended to support conditional control over generation characteristics, such as sentiment, thereby offering fine-grained steering of LLM outputs. Extensive experiments on various LLMs—including Qwen2.5-1.5B, Llama2-7B, and Llama3-8B—demonstrate that LLMBraces outperforms baseline approaches in both fine-tuning and zero-shot settings while requiring significantly fewer tunable parameters, up to 75% fewer compared to LoRA. Furthermore, LLMBraces excels in sentiment-controlled generation and toxicity reduction, highlighting its potential for flexible, controlled text generation across applications.
@@
Ying Shen Computer Science Department University of Illinois Urbana-Champaign Champaign ying22@illinois.edu                        Lifu Huang Computer Science Department UC Davis lfuhuang@ucdavis.edu
@@ 2 Related Work
Understanding LLMs Internal Working
@@ 4.1 Relevance of Sub-updates
The resulting vector ${\bm{r}}^{\ell}_{i}=[{\bm{r}}^{\ell}_{i,0},\dots,{\bm{r}}^{\ell}_{i,d_{m}}]^{\top}$ comprises individual relevance scores, with each component ${\bm{r}}^{\ell}_{i,0}$ corresponds to the $j$-th sub-update $(w^{\ell}_{i,j}\cdot{\bm{v}}^{\ell}_{j})$. In effect, the relevance score measures how well each sub-update aligns with the input context, allowing us to modulate the contribution of sub-updates based on their relevance to improve model performance.
@@ 6.1 Supervised Fine-tuning Performance
Our experiments show that As shown in Table 1, for supervised fine-tuning performance on commonsense reasoning tasks, LLMBraces achieves competitive, and even superior, performance while requiring significantly fewer trainable parameters. For instance, on the Llama3-8B model, LLMBraces outperforms LoRA by 5.2% while reducing the number of trainable parameters by approximately 38%. These results demonstrate that LLMBraces can efficiently enhance the model performance and reliability, while matching or surpassing the strong model fine-tuning based baselines, with greater parameter efficiency.
@@ 6.3 Conditional Generation
We further evaluate LLMBraces’s performance in zero-shot conditional generation. For this, we construct a new dataset, StyleDataset, by selecting 10K sentences from OpenWebText (Gokaslan and Cohen, 2019). For each sentence, we prompt Llama3.1-8B-instruct (Dubey et al., 2024) to generate a short list of descriptive attribute words that capture the sentence’s style, resulting in a dataset of 10K sentence-attribute pairs. The prompting instructions are detailed in § C.2. We then train LLMBraces on StyleDataset by pairing the attribute tokens with a positive steering value ($s=1$) and maximizing the likelihood of the corresponding sentence. Since LLMBraces is not fine-tuned specifically for sentiment or toxicity tasks, we can evaluate its zero-shot generalization on both tasks. As shown in Tables 3 and 5, LLMBraces consistently outperforms the zero-shot baseline in both sentiment control and toxic suppression. Notably, for toxic suppression, the zero-shot performance of LLMBraces even exceeds that of its fine-tuned counterpart. These results highlight LLMBraces’s robustness and adaptability in zero-shot conditional generation tasks, eliminating the need for task-specific fine-tuning.
6.4 Effect of Relevance Module
@@ 7 Conclusion
While LLMBraces is designed to be compatible with any architecture that includes a feed-forward layer, our current study focuses solely on text-based LLMs, without evaluating its applicability to multimodal language model and image generation models. In future work, we plan to extend LLMBraces to a broader range of architectures, including multimodal and image generation models, to assess its versatility and impact across diverse. Additionally, although our approach leverages sub-updates to refine model predictions, our work does not fully interpret the internal mechanisms of LLMs. Demystifying the internal states of these models remains a challenging task, and further research is needed.

### v2

@@ Abstract
Recent findings reveal that much of the knowledge in a Transformer-based Large Language Model (LLM) is encoded in its feed-forward (FFN) layers, where each FFN layer can be interpreted as the summation of sub-updates, each corresponding to a weighted column vector from the FFN’s value parameter matrix that often encodes human-interpretable concepts. In light of this, we hypothesize that model performance and behaviors can be further enhanced and controlled by modulating the contributions of these sub-updates based on their relevance to the input or target output style, and propose LLMBraces, a novel and efficient method that computes relevance scores associated with value vectors in FFN layers and leverages these scores to dynamically adjust the contribution of sub-updates. By optimizing sub-update contributions, LLMBraces refines the prediction process, leading to more accurate and reliable outputs, much like a ‘brace’ providing support and stability. Moreover, LLMBraces can be extended to support conditional control over generation characteristics, such as sentiment, thereby offering fine-grained steering of LLM outputs. Extensive experiments on various LLMs—including Qwen2.5-1.5B, Llama2-7B, and Llama3-8B—demonstrate that LLMBraces outperforms baseline approaches in both fine-tuning and zero-shot settings while requiring significantly fewer tunable parameters, up to 75% fewer compared to LoRA. Furthermore, LLMBraces excels in sentiment-controlled generation and toxicity reduction, highlighting its potential for flexible, controlled text generation across applications.
@@
Ying Shen Computer Science Department University of Illinois Urbana-Champaign ying22@illinois.edu                        Lifu Huang Computer Science Department UC Davis lfuhuang@ucdavis.edu
@@ 2 Related Work
Understanding LLMs’ Internal Working
@@ 4.1 Relevance of Sub-updates
The resulting vector ${\bm{r}}^{\ell}_{i}=[{\bm{r}}^{\ell}_{i,0},\dots,{\bm{r}}^{\ell}_{i,d_{m}}]^{\top}$ comprises individual relevance scores, with each component ${\bm{r}}^{\ell}_{i,j}$ corresponds to the $j$-th sub-update $(w^{\ell}_{i,j}\cdot{\bm{v}}^{\ell}_{j})$. In effect, the relevance score measures how well each sub-update aligns with the input context, allowing us to modulate the contribution of sub-updates based on their relevance to improve model performance.
@@ 6.1 Supervised Fine-tuning Performance
As shown in Table 1, for supervised fine-tuning performance on commonsense reasoning tasks, LLMBraces achieves competitive, and even superior, performance while requiring significantly fewer trainable parameters. For instance, on the Llama3-8B model, LLMBraces outperforms LoRA by 5.2% while reducing the number of trainable parameters by approximately 38%. These results demonstrate that LLMBraces can efficiently enhance the model performance and reliability, while matching or surpassing the strong model fine-tuning based baselines, with greater parameter efficiency.
@@ 6.3 Conditional Generation
We further evaluate LLMBraces’s performance in zero-shot conditional generation. For this, we construct a new dataset, StyleDataset, by selecting 10K sentences from OpenWebText (Gokaslan and Cohen, 2019). For each sentence, we prompt Llama3.1-8B-instruct (Dubey et al., 2024) to generate a short list of descriptive attribute words that capture the sentence’s style, resulting in a dataset of 10K sentence-attribute pairs. The prompting instructions are detailed in § C.2. We then train LLMBraces on StyleDataset by pairing the attribute tokens with a positive steering value ($s=1$) and maximizing the likelihood of the corresponding sentence. Since LLMBraces is not fine-tuned specifically for sentiment or toxicity tasks, we can evaluate its zero-shot generalization on both tasks. As shown in Tables 3 and 5, LLMBraces consistently outperforms the zero-shot baseline in both sentiment control and toxicity suppression. Notably, for toxic suppression, the zero-shot performance of LLMBraces even exceeds that of its fine-tuned counterpart. These results highlight LLMBraces’s robustness and adaptability in zero-shot conditional generation tasks, eliminating the need for task-specific fine-tuning.
@@ 7 Conclusion
While LLMBraces is designed to be compatible with any architecture that includes a feed-forward layer, our current study focuses solely on text-based LLMs, without evaluating its applicability to multimodal language model and image generation models. In future work, we plan to extend LLMBraces to a broader range of architectures, including multimodal and image generation models, to assess its versatility and impact across diverse architectures and domains. Additionally, although our approach leverages sub-updates to refine model predictions, our work does not fully interpret the internal mechanisms of LLMs. Demystifying the internal states of these models remains a challenging task, and further research is needed.
Acknowledgements
This research is supported by the award No. 2238940 from the Faculty Early Career Development Program (CAREER) of the National Science Foundation (NSF). The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.